OpenStudy (anonymous):

Can someone show me step by step to solve this word problem please!!!!!! The sum of a two-digit number and its units digit is 20, and the sum of the number and its tens digit is 16. find the number.

OpenStudy (anonymous):

we suppose that number is x and its units digit is b and its tens digit is a we have x+b=20 or 10a+b+b=20 or 10a+2b=20 x+a=16 or 10a+b+a=16 or 11a+b=16 solving the system of equations above tells a=1 and b=5 so the number is 15

OpenStudy (anonymous):

oh ok i see what you did

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

your welcome